OpenStudy (anonymous):

wat x+7>-15 i need to show how i got the answer plz help me out

OpenStudy (anonymous):

\[x > -21\] \[x \in \left( -21,\infty \right) \]

OpenStudy (anonymous):

\[\frac{ 5 }{ 7b } - \frac{ 3 }{ 2b } = \frac{ -55 }{ 42 }\] \[\frac{ 10b-21b }{ 14b^2 } = \frac{ -55 }{ 42 }\] multiplying by 14 we get \[10 - 21 = -55b/3\] -11 = -55b/3 -33/-55 = b b = 3/5